# Open Your Ears to Take a Look: A State-of-the-Art Report on the Integration of Sonification and Visualization

### v1

@@ 1.3 How to Use This Survey
With this STAR we intend to provide an overview of an emerging research field, as well as connect two mostly disjoint research communities. We hope to reach researchers from both communities, inspiring them for intertwining sonification and visualization in their future research. We see several ways of using this STAR:
@@
This STAR will be structured as follows: section 2 describes the methodology used to search and filter the literature identified as potentially relevant. In section 3 we describe our classification system and use it to discuss the survey literature. In section 4 we apply a meta perspective on the survey data, describing correlations between individual tags, as well as the co-author network of the field. In section 5 we introduce the three adjacent topics of accessibility, monitoring, and arts, which are related to our STAR, but were not systematically studied. Finally, in section 6, we offer a concluding discussion focusing on future work.
@@ 2 Method
In this section, we discuss our inclusion and exclusion criteria and the methods we used to search for the relevant literature. We used a five-stage pipeline to construct a corpus of research works that are at the intersection of visualization and sonification for data exploration and presentation.
@@ 2.1 Scope of the Surveyed Literature
Sonification and visualization share the aim of making data interpretable to their users and observers. With this shared goal, combinations of the two can be designed for numerous possible applications and contexts. Our research interest in this STAR is the combination of sonification and visualization in the context of data analysis, covering both data exploration and presentation. A work relevant for our STAR must include both, visualization and sonification of data. Therefore, a sonification with a visual interface that does not represent data is not enough to be considered relevant, neither is a visualizations with sounds that do not represent data. The work must be an academic paper, published between the years 2011 and 2023 and must be peer reviewed to be considered in our STAR.
Thinking more broadly about the combination of sonification and visualization, three additional areas of application come to mind: (1) accessibility, (2) real-time monitoring, and (3) arts. All three areas are vast, and a detailed classification of them is beyond the scope of our STAR. However, we find them relevant and inspirational for our field. Thus, we provide a brief introduction to the fields of accessibility, monitoring, and artistic contributions in section 5. In the context of accessibility, sonification can be used to support the collaboration between sighted and non-sighted individuals by mapping data to both a visual and an auditory display. In the same manner, such a design could support the collaboration between deaf individuals and individuals without hearing loss. Nevertheless, our research interest is the combination of sonification and visualization for the integrated analysis of data. Therefore, in our STAR, we consider only designs intended to be used with both the visual and the auditory senses fully available to a user. The application of real-time monitoring such as medical monitoring, monitoring of critical infrastructure, alarms, or real-time feedback of body movement, is vast and distinct from the purpose of data exploration. Especially with respect to sonification and auditory display, the field is well researched (e.g., [KIK19, SJMT19, VRGM20, $*{}^{*}$17, HRM15]), and we will not cover such designs in this STAR. Artistic contributions have the potential to be highly inspirational for our field, but require a different search method and most likely a different system of classification. Again, we decided to not systematically cover artistic contributions in this STAR, but to provide a subsection, discussing a list of representative works, that serves as a starting point for future research interests.
@@ 2.2 Search Strategy and Filtering
We base our corpus of literature on publications that the authors have already been aware of from their previous work in this field and by an extensive online literature search. The online search was a keyword-based search in the digital libraries of IEEE Xplore, ACM Digital Library, and Springer Link, which include work published at IEEE VIS, CHI, and other VGTC- and SIGCHI-sponsored venues. Furthermore, we searched the digital libraries of Eurographics, ICAD, ISon, Organised Sound, and the Sound and Music Computing Community. Figure 1 provides an overview of the different stages we used to systematically filter our database for relevant publications. To keep track of the progress of identifying relevant papers, we used a Google sheet document. The final database, including all papers and all tags by the authors, is provided in the supplemental material as a CSV file.
@@
Stage 1 – The search query we used in stage 1 was the following: ("Visualization" OR "Visualisation" OR "Visual Analytics") AND ("Sonification" OR "Auditory Display") AND NOT (centrifug* OR lys* OR homogeniz*) It consists of an AND combination of a visualization term and a sonification term combined with the exclusion of specific word-beginnings. The reason to exclude papers that hold words starting with centrifug*, lys*, and homogeniz* is the fact that sonification, also called sonication, is a term used also in biology to describe a process where sound is used to agitate particles in a sample. As papers using sonification in this context are not relevant to this STAR, we identified terms, including the three mentioned above, that are often associated with this meaning of sonification. The search in the Springer Link database especially, was sensible to the exclusion of these terms. As there could be false negatives using the exclusion of the three terms, we conducted a manual review of the paper titles that were excluded due to this strategy and restored five papers that were potentially relevant. This online search, combined with papers we were already aware of from our previous work in the field, resulted in a database holding 1498 papers. To download the respective papers and to make them available to all co-authors we used the literature management software Zotero.
@@
Stage 2 – In the second phase of our literature search, each paper title was read by two of the authors and classified into potentially relevant or irrelevant. We agreed to use an inclusive mindset for this stage, i.e., we tagged vague titles mostly as potentially relevant to not overlook too many papers at this early stage. For papers that were tagged differently by two people, the two people came to an agreement or the paper was taken to the next stage. Stage 2 resulted in a database holding 500 papers.
@@
Stage 3 – In the third phase of our literature search, each abstract was read by two of the authors and classified into potentially relevant or irrelevant. Again, we used an inclusive mindset and tagged vague abstract as potentially relevant. For papers that were tagged differently by two people, the two people came to an agreement or the paper was taken to the next stage. Stage 3 resulted in a database holding 163 papers.
@@
Stage 4 – In the fourth phase of our literature search, each paper was read by one of the authors and classified as relevant or irrelevant. As the authors had a solid common understanding of the inclusion and exclusion criteria by this stage, and as a person had the full paper information available to take a decision, one person decided on the relevance at this stage. The papers identified as relevant in stage 4 were classified using the tags explained in detail in Table 2. Each paper was first classified by one of the co-authors and their tags were later verified by a second co-author. Whenever two co-authors initially disagreed on a specific tag, they came to an agreement or the first author took a decision. Stage 4 resulted in a database holding 47 papers.
@@
Stage 5 – Furthermore, we extended the corpus of relevant papers by snowballing [Woh14], checking all incoming outgoing references of the articles matching our inclusion criteria. Snowballing was done by one of the co-authors with an exclusive mindset towards the paper titles, meaning that a vague title was not considered as relevant. Stage 5 resulted in our database holding 57 papers overall, adding 10 papers to the prior stage. During the final two stages, whenever we identified an audiovisual idiom that was published in more than one paper, we retained the most extensive version in our STAR. We identified two such cases where a design was previously published in a short paper but later expanded [YH18, MAFP19b].
@@
Figure 1: The literature search was conducted in five stages ranging from an online keyword search to using the snowballing technique on the as relevant identified papers.
@@
Paper Theme Demo Link Purpose Redundancy Visualization Idiom Soni. Technique Goal
@@
[$*{}^{*}$18] Domain Agnostic Display $\bullet$ $\bullet$ convex hull, other, bar chart $\bullet$ $\bullet$ $\bullet$
[$*{}^{*}$22a] Domain Agnostic Display [<—>] $\bullet$ $\bullet$ $\bullet$ $\bullet$ scatter plot $\bullet$ $\bullet$
@@
[$*{}^{*}$18] Earth Science $\bullet$ $\bullet$ map $\bullet$ $\bullet$
@@
[$*{}^{*}$12] Earth Science [<—>] $\bullet$ $\bullet$ map $\bullet$ $\bullet$
[$*{}^{*}$22] Earth Science [<—>] $\bullet$ $\bullet$ dot map $\bullet$ $\bullet$
@@
[$*{}^{*}$16] Medicine and Health $\bullet$ $\bullet$ slicing, volume rendering $\bullet$ $\bullet$
@@
[$*{}^{*}$23] Medicine and Health [<—>] $\bullet$ $\bullet$ map $\bullet$ $\bullet$
[$*{}^{*}$18] Medicine and Health [<—>] $\bullet$ $\bullet$ volume rendering $\bullet$ $\bullet$
@@
[$*{}^{*}$21] Medicine and Health [<—>] $\bullet$ $\bullet$ volume rendering $\bullet$ $\bullet$
[$*{}^{*}$18] Molecular Science [<—>] $\bullet$ $\bullet$ 3D molecule rendering $\bullet$ $\bullet$
@@
[$*{}^{*}$15] Molecular Science $\bullet$ $\bullet$ volume rendering $\bullet$ $\bullet$ $\bullet$
[FN18] Others $\bullet$ $\bullet$ line chart $\bullet$ $\bullet$
@@
[$*{}^{*}$12] Others [<—>] $\bullet$ $\bullet$ volume rendering $\bullet$ $\bullet$
@@ 3 Categorization and Results
In this section, we will discuss the relevant literature in detail from the perspectives of our classification. The systematic integration of sonification and visualization is a wide and diverse research field that is difficult to classify using only a handful of categories. Therefore, we decided to apply an extensive list of tags to the literature to be able to present diverse perspectives on the field, mostly concerning basic research, i.e., the basic principles that distinguishes the designs/idioms from each other.
To give readers an initial thematic overview of the field, we will start by briefly describing each of the 57 selected papers in subsection 3.1. We will do so by clustering the literature to the following topics: astronomy, medicine and health, molecular science, earth science, domain agnostic data displays, and other topics. We will then continue with a discussion on the purpose that an idiom can be designed for (subsection 3.2), followed by an analysis of idiom design possibilities (subsection 3.3). We will study several technical perspectives that were tagged individually for the sonification and the visualization aspects of each paper: the reading levels (subsection 3.4) suggested by Bertin [Ber83], the search levels (subsection 3.5) suggested by Munzner [Mun15], as well as the dataset types [Mun15] including the levels of measurement of the displayed data (subsection 3.6). We will then review different evaluation methods (subsection 3.8), different target platforms, and various possibilities of interacting with audiovisual display idioms (subsection 3.9). Finally, we study the diverse user groups and the possible goals of designers (subsection 3.10).
All the above categories and subcategories are concisely presented in two tables. Table 1 shows an overview of all papers and a subset of their most relevant tags. Table 2 provides a detailed description for each category and their respective subcategories. We also report on the total number of papers within each subcategory. The four technical categories mentioned above have distinct visualization and sonification tags which are represented in the left and right boxes in the “Num.” column, respectively. The color saturation of the boxes encodes the total number of papers in each subcategory.
@@ 3.1 Thematic Corpus Overview
Scanning our database, we identified seven astronomy related papers. Riber [Rib18] presented a prototypical virtual and interactive audio synthesizer called Planethesizer that enables users to design sonifications, especially focused on planetary data. Sonifigrapher [Rib19] is a virtual synthesizer that lets users sonify the light curves data from NASA’s exoplanet archive. Also the recently presented Sonified Hertzsprung-Russel Diagram [HPDW23] sonifies the light curves, with the diagram acting as both the visualization and the interface to choose a star to be sonified. With this design, hearing a constant pitch will inform a user about the rotation of a star. The rotation, temperature, and other parameters of planets in our solar system were also sonified by Elmquist et al. [EEBR21] in OpenSpace Sonification. Their design can be used both with conventional computer desktop environments as well as in planetarium settings, tailored towards public outreach and science communication. Public outreach is also the core of the publication Audio Universe by Harrison et al. [HTHB22]. The publication describes the design of a 35-minute audiovisual show about the the solar system integrating visualization and sonification, as well as an audiovisual animation displaying the stars in the order their appear to our eyes during dusk. Similarly, Russo and Santaguida [RS22] collaborated with NASA, celebrating the discovery of the 5000th exoplanet. Their design displays the exoplanets as they were discovered over the years. Recently, Traver presented another audiovisual installation where users can control the auditory representation of the planets using a Midi controller [TB23].
@@
We identified six medicine and health related topics in our database, out of which three are related to brain scans [GR11, RFM13, $*{}^{*}$16], two are audiovisually displaying blow flow and aneurysm models [$*{}^{*}$18, $*{}^{*}$21], and one is concerned with Covid-19 data [$*{}^{*}$23].
@@
Our database holds four idioms that we relate to molecular science. In an idiom presented by Rau et al. [$*{}^{*}$15], scientific visualization of a molecular simulation is enhanced using parameter mapping sonification and auditory icons. Among other things, their design guides the attention of a user towards visually occluded phenomena using sonification. Ballweg et al. [BBV16] use sonification with the intention of supporting chemists and structural biologists with drug design. For their sonification plug-in, they focused on tasks that were not well supported visually in a software for the interactive visualization of molecular structures called “UCSF Chimera.” In the context of biomolecules simulation, Arbon et al. [$*{}^{*}$18] developed a sonification displaying characteristics of the “free energy landscape,” a map used to study the properties of biomolecular systems. Their technique allows a user to visually inspect the physical configuration of a biomolecule while listening to their corresponding free energy landscape. Exploring the possibilities of 3D sound, Bouchara and Mones [BM20] suggested a work in progress immersive sonification model to study protein surfaces.
@@
The earth science cluster in our survey data holds eleven publications covering topics that range from oceanographic data [SAR22], wildfires [HK22], hurricanes [Bal15], and climate change [Bea11], to sonophenology [$*{}^{*}$12], seismology [HCTP14, WW15, MMU16, $*{}^{*}$22], and geospatial data displays [BF12, $*{}^{*}$18].
@@
The category of domain agnostic data display idioms in our survey data holds 14 papers. These idioms are not designed to support users from a specific domain, but are implementations tackling problems across multiple domains. Six of the papers describe software frameworks that are intended to help people design sonifications along with visual representations of their data [PC19, PCB23, LF21, DLVDCG22, CWM21, KLTW17]. Their unifying core goal is the democratization of sonification as a technique to represent data, hence making it accessible to more people, both professionals as well as domain experts. Other publications focus on basic research combining different sonification techniques such as parameter mapping sonification or model-based sonification with basic information visualizations such as scatterplots [$*{}^{*}$22a, RJ16, YH17], parallel coordinates plots [BTB23, RJ16], or line charts [FBC12] to study the potentials of audiovisual display idioms. Groppe et al.  [GKW21] studied network visualization and sonification through their design, while Malikova et al. [MAFP19b] show the potential of sonification to help users identify smallest symmetry differences in scalar fields visualizations. While most studies focus on metrics such as precision, error rates, or task completion times, the study by Du et al. [$*{}^{*}$18] explicitly investigates the sonification’s influence on user engagement.
Finally, we want to provide an overview of the remaining other 15 publications not part of the above thematic clusters. The topics related to the publications tagged as "other" are diverse, ranging from a multimodal system for analyzing business process execution data [HAR16] to an audiovisual representation of the Portuguese consumption patterns [MMM18] to a multimodal implementation of “Game of Life” using an idiom that sonifies three-dimensional cellular automata, enhancing perceived quality, patterns, and immersiveness [KAV21]. A musical sonification aimed at conveying information about running data and emotion [Rön21], an investigation into the use of timbre in sonification reveal its impact on trend identification in sonified line charts [FN18]. North et al. [NSC16] propose a hybrid version control system that employs audiovisual feedback to represent software development data such as timelines and conflicts. Alonso et al. [$*{}^{*}$12] develop an interface for product design that introduces a multisensory approach for communicating virtual objects’ geometrical data, i.e., curvature and shape. While Adhitya and Kuuskankare [AK11] propose a prototype that offers a sonification-based approach to urban design planning, several publications focus on audiovisual data representation in VR. Chabot and Braasch [CB17] present a multimodal approach to a complex stock market dataset, performed in an immersive virtual reality setting and utilizing dynamic spatialization techniques for a collaborative experience. Papachristodoulou et al. [PBM15] integrate sonification into exploring a complex neuroscience dataset using an immersive VR application, revealing the potential for improved user performance in brain region identification across different spatial resolutions. Berger and Bill[BB19] present an immersive, multisensory, and audiovisual VR application to support decision-makers and citizens about urban traffic noise issues, safety, and quality of life. An exciting approach to neuroscience through a complex dataset in immersive and interactive 3D visualization is given by Papachristodoulou et al. [PBV14], demonstrating that sonification enhances subjects’ structural understanding of the dataset, providing an additional layer of comprehension for hidden patterns in large datasets. While Joliat et al. [JMP13] combines spatialized sonification with 3D data visualization for sensor data communication, enabling data exploration with privacy protection and time-compression algorithms. Two publications focus on Explainable Artificial Intelligence (XAI). Lyu et al. [LLW21] present an interactive, audiovisual VR representation of Neural Networks, while Herrmann [Her20] designs a system that visualizes and sonifies the inner workings of a music-trained neural network, employing contrastive predictive coding. It enables the rendering of how music sounds to the artificial ear and dynamic visualization of neuron activations using a force graph layout technique.
@@
As we have now briefly touched on all 57 papers that are part of our database, we can focus on the meta level classification in the following subsections. With these more high level descriptors, we intend to provide a number of versatile perspectives on the literature. They will help us identify research gaps and opportunities for the systematic integration of sonification and visualization for the future work of both research communities in section 6. Along the discussion of these perspectives, we will present a selection of papers that are representative examples for the respective category. By doing so, we intend to provide the reader with (1) a broad exploration of the field overall, and (2) insights into the content of the actual papers themselves.
@@
Purpose
  29
Exploration
designs used for data exploration
  17
Presentation
designs used for data presentation
  11
Both
designs used for both purposes above
Visualization Design
Idiom

### v2

@@ 1.4 How to Use This Survey
With this STAR, we intend to provide an overview of an emerging research field, as well as connect two mostly disjoint research communities. We hope to reach researchers from both communities, inspiring them to intertwine sonification and visualization in their future research. We see several ways of using this STAR:
@@
This STAR will be structured as follows: section 2 describes the methodology used to search and filter the literature identified as potentially relevant. In section 3, we describe our classification system and use it to discuss the survey literature. In section 4, we apply a meta-perspective on the survey data, describing correlations between individual tags, as well as the co-author network of the field. In section 5 we introduce the three adjacent topics of accessibility, monitoring, and arts, which are related to our STAR, but were not studied systematically. Finally, in section 6, we offer a concluding discussion focusing on future work.
@@ 2 Method
In this section, we discuss our inclusion and exclusion criteria and the methods we used to search for the relevant literature. We used a five-stage pipeline to construct a corpus of research that is at the intersection of visualization and sonification for data exploration and presentation (see Figure 1).
@@ 2.1 Scope of the Surveyed Literature
Sonification and visualization share the aim of making data interpretable to their users and observers. With this shared goal, combinations of the two can be designed for numerous possible applications and contexts. Our research interest in this STAR is the combination of sonification and visualization in the context of data analysis, covering both data exploration and presentation. A work that is relevant to our STAR must include both visualization and sonification of data. Therefore, a sonification with a visual interface that does not represent data is not enough to be considered relevant, and neither is a visualization with sounds that do not represent data. The work must be an academic paper published between the years 2011 and 2023 and must be peer-reviewed to be considered in our STAR.
Thinking more broadly about the combination of sonification and visualization, three additional areas of application come to mind: (1) accessibility, (2) real-time monitoring, and (3) arts. All three areas are vast, and a detailed classification of them is beyond the scope of our STAR. However, we find them relevant and inspirational for our field. Thus, we provide a brief introduction to the fields of accessibility, monitoring, and artistic contributions in section 5. In the context of accessibility, sonification can be used to support the collaboration between blind and sighted users by mapping data to both an auditory and a visual display. In the same manner, such a design could support the collaboration between deaf individuals and individuals without hearing impairment. Nevertheless, our research interest is the combination of sonification and visualization for the integrated analysis of data. Therefore, in our STAR, we consider only designs intended to be used with both the visual and the auditory senses fully available to a user. The application of real-time monitoring, such as medical monitoring, critical infrastructure monitoring, alarms, or real-time feedback on body movement, is vast and distinct from the purpose of data exploration. Especially with respect to sonification and auditory display, the field is well-researched (e.g., [KIK19, SJMT19, VRGM20, WMY17, HRM15]), and we will not cover such designs in this STAR. Artistic contributions have the potential to be highly inspirational for our field but require a different search method and, most likely, a different system of classification. Again, we decided not to systematically cover artistic contributions in this STAR but to provide a subsection discussing a list of representative works that may serve as a starting point for future research interests.
@@ 2.2 Search Strategy and Filtering
We base our corpus of literature on (1) publications that the authors have already been aware of from their previous work in this field and (2) an extensive online literature search. The online search was a keyword-based search in the digital libraries of IEEE Xplore, ACM Digital Library, and Springer Link, which include work published at IEEE VIS, CHI, and other VGTC- and SIGCHI-sponsored venues. Furthermore, we searched the digital libraries of Eurographics, ICAD, ISon, Organised Sound, and the Sound and Music Computing Community. Figure 1 provides an overview of the different stages we used to systematically filter our database for relevant publications. To keep track of the progress, we used a Google sheet document. The final database, including all papers and all tags by the authors, is provided in the supplemental material as a CSV file.
@@
Stage 1 – The search query we used in stage 1 was the following: ("Visualization" OR "Visualisation" OR "Visual Analytics") AND ("Sonification" OR "Auditory Display") AND NOT (centrifug* OR lys* OR homogeniz*). It consists of an AND combination of a visualization term and a sonification term combined with the exclusion of specific word beginnings. While the terms “visualization” and “visual analytics” are well known, our audio-related search terms are likely to be less well-known overall. Nevertheless, we decided not to include search terms such as music, sound, or tone in our online search. Searching for publications with such broad terminology would have resulted in too many papers to be scanned. Also, we argue that academic publications interested in the integration of visualization and sonification are likely using the appropriate terminology. The reason to exclude papers that hold words starting with centrifug*, lys*, and homogeniz* is the fact that sonification, also called sonication, is a term also used in biology to describe a process where sound is used to agitate particles in a sample. As papers using sonification in this context are not relevant to this STAR, we identified terms, including the three mentioned above, that are often associated with this meaning of sonification. The search in the Springer Link database was especially sensitive to the exclusion of these terms. As there could be false negatives using the exclusion of the three terms, we manually reviewed the paper titles that were excluded due to this strategy and restored five potentially relevant papers. This online search, combined with papers we were already aware of from our previous work in the field, resulted in a database holding 1498 papers. We used the literature management software Zotero to download the respective papers and to make them available to all co-authors.
@@
Stage 2 – In the second phase of our literature search, each paper title was read by two of the authors and classified into potentially relevant or irrelevant. We agreed to use an inclusive mindset for this stage, i.e., we tagged vague titles mostly as potentially relevant so as not to overlook many papers at this early stage. For papers that were tagged differently by two people, the two people came to an agreement, or the paper was taken to the next stage. Stage 2 resulted in a database holding 500 papers.
@@
Stage 3 – In the third phase of our literature search, each abstract was read by two of the authors and classified into potentially relevant or irrelevant. Again, we used an inclusive mindset and tagged vague abstracts as potentially relevant. For papers that were tagged differently by two people, the two people came to an agreement, or the paper was taken to the next stage. Stage 3 resulted in a database holding 163 papers.
@@
Stage 4 – In the fourth phase of our literature search, each paper was read by one of the authors and classified as relevant or irrelevant. As the authors had a solid common understanding of the inclusion and exclusion criteria by this stage, and as a person had the full paper information available to make a decision, one person decided on the relevance at this stage. The papers identified as relevant in stage 4 were classified using the tags explained in detail in Table 2. Each paper was first classified by one of the co-authors, and their tags were later verified by a second co-author. Whenever two co-authors initially disagreed on a specific tag, they came to an agreement, or the first author made a decision. Stage 4 resulted in a database holding 47 papers.
@@
Stage 5 – Furthermore, we extended the corpus of relevant papers by snowballing [Woh14], checking all incoming and outgoing references of the articles matching our inclusion criteria. Snowballing was done by one of the co-authors with an exclusive mindset towards the paper titles, meaning that a vague title was not considered relevant. Stage 5 resulted in our database holding 57 papers overall, adding ten papers to the prior stage. During the final two stages, whenever we identified an audiovisual idiom that was published in more than one paper, we retained the most extensive version in our STAR. We identified two such cases where a design was previously published in a short paper but later expanded [YH18, MAFP19b]. It is notable that a considerable number of papers that are part of our final scope were published at the International Conference on Auditory Display (ICAD). This is not surprising as this venue is the most recognized venue to publish sonification work holding the largest single corpus of work in the field [BB12, AG21, GVEZ23]. It is also not surprising that it is a sonification venue that is prominent, as there is a structural imbalance between the domains. Many sonification designs, in general, also include some sort of visualization in their design, but that is not true the other way around.
@@
Figure 1: The literature search was conducted in five stages, ranging from an online keyword search to using the snowballing technique on the papers identified as relevant. Figure inspired by [YM22].
@@
Paper Theme Demo Link Purpose Redundancy Visualization Idiom Sonification Technique Goal
@@
[DCM18] Domain Agnostic Display $\bullet$ $\bullet$ convex hull, other, bar chart $\bullet$ $\bullet$ $\bullet$
[ERI22] Domain Agnostic Display [<—>] $\bullet$ $\bullet$ $\bullet$ $\bullet$ scatter plot $\bullet$ $\bullet$
@@
[FN18] Domain Agnostic Display $\bullet$ $\bullet$ line chart $\bullet$ $\bullet$
@@
[GDAS18] Earth Science $\bullet$ $\bullet$ map $\bullet$ $\bullet$
@@
[NRL12] Earth Science [<—>] $\bullet$ $\bullet$ map $\bullet$ $\bullet$
[PFH22] Earth Science [<—>] $\bullet$ $\bullet$ dot map $\bullet$ $\bullet$
@@
[GRK16] Medicine and Health $\bullet$ $\bullet$ slicing, volume rendering $\bullet$ $\bullet$
@@
[LSB23] Medicine and Health [<—>] $\bullet$ $\bullet$ map $\bullet$ $\bullet$
[MNW18] Medicine and Health [<—>] $\bullet$ $\bullet$ volume rendering $\bullet$ $\bullet$
@@
[TMN21] Medicine and Health [<—>] $\bullet$ $\bullet$ volume rendering $\bullet$ $\bullet$
[AJB18] Molecular Science [<—>] $\bullet$ $\bullet$ 3D molecule rendering $\bullet$ $\bullet$
@@
[RFK15] Molecular Science $\bullet$ $\bullet$ volume rendering $\bullet$ $\bullet$ $\bullet$
@@
[ASH12] Others [<—>] $\bullet$ $\bullet$ volume rendering $\bullet$ $\bullet$
@@ 3 Categorization and Results
In this section, we will discuss the relevant literature in detail from the perspectives of our classification. The systematic integration of sonification and visualization is a wide and diverse research field that is difficult to classify using only a handful of categories. Therefore, we decided to apply an extensive list of tags to the literature to be able to present diverse perspectives on the field, mostly concerning basic research, i.e., the basic principles that distinguish the designs/idioms from each other.
To give readers an initial thematic overview of the field, we will start by briefly describing each of the 57 selected papers in subsection 3.1. We will do so by clustering the literature to the following topics: astronomy, medicine and health, molecular science, earth science, domain agnostic data displays, and other topics. We will then continue with a discussion on the purpose that an idiom can be designed for (subsection 3.2), followed by an analysis of idiom design possibilities (subsection 3.3). We will study several technical perspectives that were tagged individually for the sonification and the visualization aspects of each paper: the reading levels (subsection 3.4) suggested by Bertin [Ber83], the search levels (subsection 3.5) suggested by Munzner [Mun15], as well as the dataset types [Mun15] including the levels of measurement of the displayed data (subsection 3.6). We will then review different levels of mapping redundancy (subsection 3.7), different evaluation methods (subsection 3.8), different target platforms, and various possibilities of interacting with audiovisual display idioms (subsection 3.9). Finally, we study the diverse user groups and the possible goals of designers (subsection 3.10).
All the above categories and subcategories are concisely presented in two tables. Table 1 shows an overview of all papers and a subset of their most relevant tags. Table 2 provides a detailed description for each category and their respective subcategories. We also report on the total number of papers within each subcategory. The four technical categories mentioned above have distinct visualization and sonification tags, which are represented in the left and right boxes in the “Num.” column. The luminance of the boxes encodes the total number of papers in each subcategory.
@@ 3.1 Thematic Corpus Overview
Before we employ the classification system described in Table 2, we want to provide a thematic corpus overview. This overview is intended for readers who are interested in a special field of application, such as astronomy or earth sciences. While we do not intend to provide detailed descriptions of the individual articles at this point, we will present more detailed descriptions in the later subsections 3.2 to 3.10.
@@
Scanning our database, we identified seven astronomy related papers. Riber [Rib18] presented a prototypical virtual and interactive audio synthesizer called Planethesizer that enables its users to design sonifications, especially focused on planetary data. Sonifigrapher [Rib19] is a virtual synthesizer that lets users sonify the light curves data from NASA’s exoplanet archive. Also, the recently presented Sonified Hertzsprung-Russel Diagram [HPDW23] sonifies the light curves, with the diagram acting as both the visualization and the interface to choose a star to be sonified. With this design, hearing a constant pitch will inform a user about the rotation of a star. The rotation, temperature, and other parameters of planets in our solar system were also sonified by Elmquist et al. [EEBR21] in OpenSpace Sonification. Their design can be used both with conventional computer desktop environments and in planetarium settings, and they are tailored towards public outreach and science communication. Public outreach is also the core of the publication Audio Universe by Harrison et al. [HTHB22]. The publication describes the design of a 35-minute audiovisual show about the solar system integrating visualization and sonification, as well as an audiovisual animation displaying the stars in the same order they appear to our eyes during dusk. Similarly, Russo and Santaguida [RS22] collaborated with NASA, celebrating the discovery of the 5000th exoplanet. Their design displays the exoplanets as they were discovered over the years. Recently, Traver presented another audiovisual installation where users can control the auditory representation of the planets using a Midi controller [TB23].
@@
We identified six medicine and health related topics in our database, out of which three are related to brain scans [GR11, RFM13, GRK16], two are audiovisually displaying blow flow and aneurysm models [MNW18, TMN21], and one is concerned with Covid-19 data [LSB23].
@@
Our database holds four idioms that we relate to molecular science. In an idiom presented by Rau et al. [RFK15], scientific visualization of a molecular simulation is enhanced using parameter mapping sonification and auditory icons. Among other things, their design guides the attention of a user towards visually occluded phenomena using sonification. Ballweg et al. [BBV16] use sonification with the intention of supporting chemists and structural biologists with drug design. For their sonification plug-in, they focused on tasks that were not well supported visually in a software for the interactive visualization of molecular structures called “UCSF Chimera.” In the context of biomolecules simulation, Arbon et al. [AJB18] developed a sonification displaying characteristics of the “free energy landscape,” a map used to study the properties of biomolecular systems. Their technique allows a user to visually inspect the physical configuration of a biomolecule while listening to their corresponding free energy landscape. Exploring the possibilities of 3D sound, Bouchara and Mones [BM20] suggested a work-in-progress immersive sonification model to study protein surfaces.
@@
The earth science cluster in our survey data holds eleven publications covering topics that range from oceanographic data [SAR22], wildfires [HK22], hurricanes [Bal15], and climate change [Bea11], to sonophenology [NRL12], seismology [HCTP14, WW15, MMU16, PFH22], and geospatial data displays [BF12, GDAS18].
@@
The category of domain agnostic data display idioms in our survey data holds 15 papers. These idioms are not designed to support users from a specific domain but are implementations tackling problems across multiple domains. Six of the papers describe software frameworks that are intended to help people design sonifications along with visual representations of their data [PC19, PCB23, LF21, DLVDCG22, CWM21, KLTW17]. Their unifying core goal is the democratization of sonification as a technique to represent data, hence making it accessible to more people, both professionals as well as domain experts. Other publications focus on basic research combining different sonification techniques such as parameter mapping sonification or model-based sonification with basic information visualizations such as scatterplots [ERI22, RJ16, YH17], parallel coordinates plots [BTB23, RJ16], or line charts [FBC12, FN18] to study the potentials of audiovisual display idioms. Groppe et al.  [GKW21] studied network visualization and sonification through their design, while Malikova et al. [MAFP19b] show the potential of sonification to help users identify smallest symmetry differences in scalar fields visualizations. While most studies focus on metrics such as precision, error rates, or task completion times, the study by Du et al. [DCM18] explicitly investigates the sonification’s influence on user engagement.
Finally, we want to provide a brief overview of the remaining other 14 publications not part of the above thematic clusters. The topics related to the publications tagged as “other” are diverse, ranging from a multimodal system for analyzing business process execution data [HAR16] or Git version control data [NSC16], to an audiovisual representation of the Portuguese consumption patterns [MMM18], to a multimodal implementation of “Game of Life” [KAV21], and a musical sonification aimed at conveying information about running data and emotion [Rön21]. Several publications focus on audiovisual data representation in virtual or extended realities [CB17, PBV14, PBM15, BB19, JMP13] and two publications focus on explainable AI [LLW21, Her20]. Alonso et al. [ASH12] presented an interface for product design that communicates a virtual object’s geometrical shape using visual, haptic, and auditory stimuli, and Adhitya and Kuuskankare [AK11] proposed a prototype that offers a sonification-based approach to urban design planning.
@@
As we have now presented a thematic overview of all 57 papers that are part of our database, we can focus on the meta-level classification in the following subsections. With these more high-level descriptors, we intend to provide a number of versatile perspectives on the literature. They will help us identify research gaps and opportunities for the systematic integration of sonification and visualization for the future work of both research communities in section 6. Along the discussion of these perspectives, we will present selected papers that are representative examples for the respective category. By doing so, we intend to provide the reader with (1) a broad exploration of the field overall and (2) insights into the content of the actual papers themselves.
@@
Purpose    29 Exploration designs used for data exploration
   17 Presentation designs used for data presentation
   11 Both designs used for both purposes above